Given the following data:
Flexible budget variance 100 F
Efficiency variance 50% of spending variance, and unfavorable
Standard rate 10
Machie hours 200
Find
Standard hours

$220

$210,000

150,000

100,000
Answer #1
Statement showing Computations
Spending variance = 100*50% = 50 U
DLEV = (SH-AH)SR
-50 = (SH -200)10
-5 =(SH -200)
Standard hours = 200 - 5 =195 hours